How many grams of Chlorine will react completely with 32.9 g of Aluminum?


​   

1
Expert's answer
2021-10-28T07:48:08-0400

2Al + 3Cl2 → 2AlCl3

MM(Al) = 26.98 g/mol

n(Al) "= \frac{32.9}{26.98} = 1.22 \;mol"

According to the reaction:

n(Cl2) "= \frac{3}{2}n(Al) = \frac{3}{2} \times 1.22 = 1.83 \;mol"

MM(Cl2) = 70.9 g/mol

m(Cl2) "= 1.83 \times 70.9 = 129.75 \;g"

Answer: 129.75 g